# General Seemingly Unrelated Local Projections

### v2

@@
We provide a framework for efficiently estimating impulse response functions with Local Projections (LPs). Our approach offers a Bayesian treatment for LPs with Instrumental Variables, accommodating multiple shocks and instruments per shock, accounts for autocorrelation in multi-step forecasts by jointly modeling all LPs as a seemingly unrelated system of equations, defines a flexible yet parsimonious joint prior for impulse responses based on a Gaussian Process, allows for joint inference about the entire vector of impulse responses, and uses all available data across horizons by imputing missing values.
@@
Contact: Michael Pfarrhofer (mpfarrho@wu.ac.at), Department of Economics, WU Vienna University of Economics and Business. Codes and replication files are available at github.com/mpfarrho/sulp. We would like to thank Gary Koop and seminar participants at the Bank of Canada and the University of Strathclyde for useful comments.
@@ 1. Introduction
In finite samples, there are (bias-variance) trade-offs between the simplicity of LPs and the structure that VARs and related models impose (Li et al., 2024). [^2] Asymptotically, both LPs and VARs estimate the same IRFs (Plagborg-Møller and Wolf, 2021). While more robust against misspecification, LP estimators are typically less efficient, and the estimated impulse responses can behave erratically since horizon-specific LPs are treated independently. In such cases, some form of regularization is generally beneficial, and can be used to introduce prior information about the shape of the IRFs. Furthermore, heteroskedasticity and autocorrelation consistent (HAC) standard errors do not exploit the available information about the correlation structure of the forecast errors across horizons.
@@
In this paper, we present a general framework to tackle these issues jointly while also confronting issues that are often set aside in LPs, such as the joint estimation of impulse responses to various shocks, in which case we need to tackle the issue that instruments for different shocks may be correlated in practice, as we show in one application. In that case, researchers need to take a stand on whether the correlation stems purely from noise (and as such needs to be filtered out) or if there is a common component/shock that itself should be studied. We investigate both alternatives. In addition, we explore how estimating LPs as a system of seemingly unrelated regressions takes the correlation of forecast errors into account.
@@
This joint modeling enables us to design Bayesian machine learning priors on the joint distribution of the impulse responses. These are based on Gaussian processes (GPs) that can be used to introduce information on the shape and magnitudes of the dynamic responses. To decide on the importance of this information, we use Bayesian shrinkage and regularization techniques that researchers in macroeconomics are familiar with.
@@
Our approach builds a pseudo-likelihood that allows us to use the full Bayesian toolkit. [^3] We elaborate on why we use a pseudo-likelihood in Section 2.1. For related discussions on likelihood-based inference for LPs, see also Tanaka (2020b). As such, the construction of error bands and related objects is straightforward using the output of our posterior sampling algorithm, a tricky issue when other modes of regularization are used. Bayesian inference in high-dimensional parameter spaces can seem daunting to researchers used to the ease of OLS-based estimation of LPs. To help in that regard, we use a hierarchical model so that only very few prior hyperparameters need to be set. And for those hyperparameters, we offer benchmark values. Our approach estimates LPs on standardized data (before translating coefficients back to the original scale if desired), so the prior can be used across applications easily.
@@
There are several important contributions dealing with regularization in LPs. In a frequentist context, Barnichon and Brownlees (2019) use splines to achieve regularization. They note the aforementioned conceptual difficulty in constructing error bands in this context. Ferreira et al. (2023) instead use a Bayesian approach for LPs, but estimate the model one horizon at a time without directly taking into account the correlation of the forecast error, requiring them to ex-post adjust standard errors using a frequentist, HAC-based, approach. [^5] Brugnolini et al. (2024) build on Ferreira et al. (2023) to estimate nonlinear LPs. Tanaka (2020a) also relies on Bayesian inference, and, to define a (joint) roughness penalty prior on the impulse response vector, estimates a system of LPs as a whole. Neither of these latter papers, however, takes into account instruments, a major part of our contribution. Aruoba and Drechsel (2024) also estimate a system of local projections, but stack equations across variables, not across horizons. [^6] They mention that stacking across horizons is also possible, but do not estimate a system of local projections across variables and horizons due to the implied computational burden.
Since we jointly estimate all impulse responses across horizons (and in some cases also across various shocks), our approach automatically allows for joint inference on impulse responses, a topic that has recently received substantial attention both in the VAR (Inoue and Kilian, 2022) and LP (Inoue et al., 2024) literature. [^7] Some papers estimate impulse responses directly using moving average models where joint inference is also natural (Barnichon and Matthes, 2018; Plagborg-Møller, 2019). More generally, our approach offers an alternative to recent frequentist approaches to inference in LPs (Montiel Olea and Plagborg-Møller, 2021; Xu, 2023).
Finally, our approach is set up to deal with missing values. LPs often require researchers to drop some data or use different sample sizes for different horizons because of their structure as estimated direct forecasts. We instead impute missing values within a unified framework, allowing researchers to use the full sample across all horizons.
The remainder of the paper is structured as follows. Section 2 introduces our econometric framework whereas Section 3 offers a discussion on model extensions to showcase the generality of our approach. We illustrate the model based on synthetic and real data in Section 4 and offer a summary and conclusions in Section 5.
@@ 2.1. Local projections as seemingly unrelated regressions
This section begins by discussing the relationship between dynamic models and LPs by using a simple example. It motivates our general modeling approach, which we describe in more detail in the next section. Let $w_{t}$ denote our variable of interest, which follows an AR(1) process:
@@
$w_{t}$ $=b_{0}w_{t-1}+u_{t}^{(0)},$
[…]
$w_{t+\tilde{H}}$ $=b_{h}w_{t-1}+u_{t+\tilde{H}}^{(\tilde{H})},$
@@
All these techniques, however, are based on treating each of the forecasting equations as separate estimation problems. [^9] One exception that takes a Bayesian stance is Tanaka (2020a) who relies on a similar representation as we do, and uses it as an estimation device to elicit a joint prior on $\bm{b}$. This has implications for the efficiency of the estimators, since $b_{i}$ and $b_{j}$ are treated independently from each other for each $i\neq j$. Moreover, impulse responses in dynamic models typically feature patterns such as persistence and smoothness, implying that an independent treatment of each horizon complicates incorporating potentially available prior knowledge about the shape of the IRF.
For these (and other reasons), our approach is different. Let $\bm{y}_{t}=(w_{t},\dots,w_{t+\tilde{H}})^{\prime}$; we estimate the system of Equations (1) jointly and assume a full covariance matrix between the shocks to the different horizons:
@@
with $\bm{b}=(b_{0},\dots,b_{H})^{\prime}$ and $\bm{u}_{t}=(u^{(0)}_{t},\dots,u^{(\tilde{H})}_{t+\tilde{H}})^{\prime}$. Note that we jointly estimate the entire vector of impulse responses $\bm{b}$, allowing us to easily make probability statements about this relationship between different elements of this vector. Joint inference for impulse responses has received substantial interest — see Inoue et al. (2024) for LPs in a frequentist context and Inoue and Kilian (2022) in a Bayesian VAR context, something our approach can naturally deliver. These aspects are discussed in detail in Section 2.3.
@@
While the forecast errors of the AR process are uncorrelated, the residuals of the LPs are correlated across horizons because they represent multi-step forecast errors. Our approach takes into account the correlations across horizons — Equation (2) motivates our Bayesian approach to estimate $\bm{b}$, taking into account the correlation structure between the shocks, determined by $\tilde{\bm{Q}}$. We achieve this by introducing a pseudo-likelihood (see also Ferreira et al., 2023, who, however, do not jointly estimate the entire system of equations and instead rely on an ex-post adjustment of the posterior variance of impulse responses), which we build as follows.
@@
Second, at a fundamental level, why is Equation (5) a pseudo-likelihood and why don’t we write down the full likelihood function? The answer is that the set of LPs does not constitute a generative model and as such the prediction error decomposition cannot be applied. To see this, note that a standard prediction error decomposition would force us to use $\mathcal{N}(\bm{y}_{t}|\bm{y}^{t-1},\bm{b},\bm{\Sigma}_{u})$, where the superscript indicates information up to $t-1$. But since $\bm{y}_{t}$ and $\bm{y}_{t-1}$ have elements in common, the prediction error for those elements would be zero, while a standard application of LPs would want to exploit information in the forecast errors for all horizons and time periods. We thus view our pseudo-likelihood as a convenient device that allows us to use the Bayesian toolkit when estimating LPs. Writing down the likelihood for the entire system takes into account correlations across horizons. To see this, it is useful to decompose $\hat{p}(\bm{y}_{t}|\bm{b},\bm{\Sigma}_{u})$ into: [^11] This decomposition is a prediction-error decomposition, but for one time-period $t$ across all horizons; the problem we just discussed arises when considering a prediction-error decomposition across $t$.
@@
where ${\bm{Q}}$ is lower triangular with unit diagonal and ${\bm{\Omega}}=\text{diag}({\omega}_{1},\dots,{\omega}_{H})$ are the error variances of $\bm{e}_{t}$, the uncorrelated errors recovered by the Cholesky decomposition, see also Tanaka (2020b). Note that each horizon has a different variance $\omega_{j}$ and hence we have heteroskedasticity across the different forecast horizons. This is in contrast to Equation (3). We view this increased flexibility as an advantage of our approach, even though it is not strictly needed in this specific application, as a comparison of Equations (2) and the expressions we turn to next show.
@@
$\bm{u}_{t}=\bm{Q}\bm{e}_{t},\quad\bm{e}_{t}\sim\mathcal{N}(\bm{0},\bm{\Omega}),\quad\bm{e}_{t}=(e_{t}^{(0)},e_{t+1}^{(1)},\ldots,e_{t+\tilde{H}}^{(\tilde{H})})^{\prime},$
@@
Hence, the $h$-step-ahead LP forecast errors are a function of the independent errors across the preceding $h-1$ horizons as in Equation (1). Our pseudo-likelihood controls for $e_{t}^{(j)}$ for all $j<h$ in the LP for horizon $h$, as these are a function of the parameters of the model and the relevant lags of $w_{t}$. Throughout, we assume that the horizon 00 forecast error $u_{t}^{(0)}$ is serially uncorrelated, a standard assumption in the literature.
To gain a better understanding of the relationship between the different errors, the model can be written in full data notation. When $e_{t+h}^{(i)}=e_{t+h}^{(j)}=e_{t+h}$, we obtain: [^12] A variant capable of strictly enforcing this is to use a state space representation and treat the one-step-ahead prediction errors as unobserved states. We provide details in Appendix A.
@@
where the $t$th row of $\bm{U}$ is $\bm{u}_{t}^{\prime}$. The $j$th (for $j=h+1$) column of $\bm{U}$, denoted $\bm{U}_{\bullet j}$, can be stated as:
@@
[…]
@@ 2.2. General seemingly unrelated linear projections
We start by generalizing the ideas laid out in the previous section. Our goal is to estimate the dynamic response of a target variable $w_{t}$ to a change in a scalar time series $x_{t}$ conditional on a possibly large panel of additional variables stored in $\bm{r}_{t}$ and $\bm{s}_{t}$. [^13] This framework in principle enables us to compute impulse responses for any number of variables, by simply replacing the dependent variable $w_{t}$ in Equation (7) with a vector of dependent variables $\bm{w}_{t}=(w_{1t},\ldots,w_{Mt})^{\prime}$. These are of dimension $n_{r}$ and $n_{s}$, respectively. The vectors $\bm{r}_{t}$ and $\bm{s}_{t}$ are used to distinguish between predetermined and simultaneously determined covariates. Let $\bm{d}_{t}=(w_{t},x_{t},\bm{r}_{t}^{\prime},\bm{s}_{t}^{\prime})^{\prime}$ and define $\bm{z}_{t}=(\bm{r}_{t}^{\prime},\bm{d}_{t-1}^{\prime},\ldots,\bm{d}_{t-P}^{\prime})^{\prime}$ as a vector of contemporaneous and lagged controls with associated $k\times 1$ parameter vector $\bm{\gamma}_{h}$ with $k=n_{r}+P(n_{r}+n_{s})$.
@@
where $\bm{\beta}=(\beta_{0},\beta_{1},\ldots,\beta_{\tilde{H}})^{\prime}$ is our main object of interest, $\bm{Z}_{t}=(\bm{I}_{H}\otimes\bm{z}_{t}^{\prime})$ is of size $H\times kH$ and $\bm{\gamma}=(\bm{\gamma}_{0}^{\prime},\bm{\gamma}_{1}^{\prime},\ldots,\bm{\gamma}_{\tilde{H}}^{\prime})^{\prime}$ includes the coefficients associated with the controls across horizons.

### v3

@@
We develop a Bayesian framework for the efficient estimation of impulse responses using Local Projections (LPs) with instrumental variables. It accommodates multiple shocks and instruments, accounts for autocorrelation in multi-step forecasts by jointly modeling all LPs as a seemingly unrelated system of equations, defines a flexible yet parsimonious joint prior for impulse responses based on a Gaussian Process, and allows for joint inference about the entire vector of impulse responses. We show via Monte Carlo simulations that our approach delivers more accurate point and uncertainty estimates than standard methods. To address potential misspecification, we propose an optional robustification step based on power posteriors.
@@
Contact: Michael Pfarrhofer (mpfarrho@wu.ac.at), Department of Economics, WU Vienna University of Economics and Business. We would like to thank Gary Koop and seminar participants at the Bank of Canada, the University of Strathclyde and the 15th RCEA Bayesian Econometrics Workshop for useful comments.
@@ 1. Introduction
In finite samples, there is a (bias–variance) trade-off between the simplicity of LPs and the structure that VARs and related models impose (Li et al., 2024). [^2] Asymptotically, both LPs and VARs estimate the same IRFs (Plagborg-Møller and Wolf, 2021). Additional discussions on the relationship between VAR-based and LP-based inference about IRFs are provided in Ludwig (2024) and Baumeister (2025). While more robust against misspecification, LP estimators are typically less efficient, and the estimated impulse responses can behave erratically since horizon-specific LPs are treated independently. In such cases, some form of regularization is generally beneficial and can be used to introduce prior information about the shape of the IRFs. Furthermore, heteroskedasticity and autocorrelation consistent (HAC) standard errors do not exploit the available information about the correlation structure of the forecast errors across horizons.
@@
In this paper, we present a general framework to tackle these issues jointly while also confronting issues that are often set aside in LPs, such as the joint estimation of impulse responses to various shocks, in which case we need to tackle the issue that instruments for different shocks may be correlated in practice, as we show in one application. In that case, researchers need to take a stand on whether the correlation stems from noise (and needs to be filtered out) or if there is a common component/shock that itself should be studied. We investigate both alternatives. In addition, we explore how estimating LPs as a system of seemingly unrelated regressions takes the correlation of forecast errors into account.
@@
This joint modeling enables us to design Bayesian machine learning priors on the joint distribution of the impulse responses. These are based on Gaussian processes (GPs), a nonparametric method popular in statistics and computer science, that can be used to introduce information on the shape and magnitudes of the dynamic responses. To decide on the importance of this information, we use Bayesian shrinkage and regularization techniques that researchers in macroeconomics are familiar with.
Our approach constructs a pseudo-likelihood that allows us to apply the full Bayesian toolkit. [^3] We explain the use of a pseudo-likelihood in Section 2.1. This setup makes it straightforward to construct error bands and related objects using the output of our posterior sampler-an area where other regularization methods often face complications. While relying on a pseudo-likelihood introduces the potential for misspecification, we address this concern in two ways. First, we estimate the local projections jointly as a system and impose a flexible nonparametric prior. As we show in Section 4, this leads to improved coverage probabilities relative to standard local projection specifications. Second, to further guard against residual misspecification, we introduce a novel, optional post-processing step that aligns the posterior with a power posterior (see, e.g., Holmes and Walker, 2017; Grünwald and Van Ommen, 2017). This adjustment offers robustness to a wide range of misspecifications and inherits desirable theoretical properties (Bhattacharya et al., 2019).
@@
Bayesian inference in high-dimensional parameter spaces can seem daunting to researchers used to the ease of OLS-based estimation of LPs. To help in that regard, we use a hierarchical model so that only very few prior hyperparameters need to be set. And for those hyperparameters, we offer benchmark values. Our approach estimates LPs on standardized data (before translating coefficients back to the original scale if desired), so the prior can be used across applications easily.
@@
There are several important contributions dealing with regularization in LPs. In a frequentist context, Barnichon and Brownlees (2019) use splines to achieve regularization. They note the aforementioned conceptual difficulty in constructing error bands in this context. Ferreira et al. (2025) instead use a Bayesian approach for LPs, but estimate the model one horizon at a time without directly taking into account the correlation of the forecast error, requiring them to ex-post adjust standard errors using a frequentist, HAC-based, approach. Tanaka (2020) also relies on Bayesian inference, and, to define a (joint) roughness penalty prior on the impulse response vector, estimates a system of LPs as a whole. Neither of these latter papers, however, takes into account instruments, a major part of our contribution. Aruoba and Drechsel (2024) also estimate a system of local projections, but stack equations across variables, not across horizons. [^5] They mention that stacking across horizons is also possible, but do not estimate a system of local projections across variables and horizons due to the implied computational burden.
Since we jointly estimate all impulse responses across horizons (and in some cases also across various shocks), our approach automatically allows for joint inference on impulse responses, a topic that has recently received substantial attention both in the VAR (Inoue and Kilian, 2022) and LP (Inoue et al., 2025) literature. [^6] Some papers estimate impulse responses directly using moving average models where joint inference is also natural (Barnichon and Matthes, 2018; Plagborg-Møller, 2019). More generally, our approach offers an alternative to recent frequentist approaches to inference in LPs (Montiel Olea and Plagborg-Møller, 2021; Xu, 2023).
Finally, our framework is set up to deal with missing values. LPs often require researchers to drop some data or use different sample sizes for different horizons because of their structure as estimated direct forecasts. We instead impute missing values within a unified framework, allowing researchers to use the full sample across all horizons.
The remainder of the paper is structured as follows. Section 2 introduces our econometric framework, whereas Section 3 offers a discussion on model extensions to showcase the generality of our approach. We illustrate the model based on synthetic and real data in Sections 4 and 5 and offer a summary and conclusions in Section 6.
@@ 2.1. Local projections as seemingly unrelated regressions
This section begins by discussing the relationship between dynamic models and LPs using a simple example. It motivates our general modeling approach, which we describe in more detail in the next section. Let $w_{t}$ denote our variable of interest, which follows an AR(1) process:
@@
$w_{t}=b_{0}w_{t-1}+u_{t}^{(0)}\quad\ldots\quad w_{t+\tilde{H}}=b_{h}w_{t-1}+u_{t+\tilde{H}}^{(\tilde{H})}$, (1)
@@
These are LPs in the spirit of Jordà (2005). However, doing so by OLS neglects the autocorrelation in forecast errors, which follow a moving average process of order $h$. As mentioned above, Lusompa (2023) shows that the MA($h$) structure of the forecast error can be obtained under mild regularity conditions and does not rely on the data-generating process (DGP) being an ARMA process of finite order as long as one assumes stationarity, which we do not have to do with our direct assumption of an AR process. [^7] Note that different to the notation of Lusompa (2023), in our case $h=0$ in Equation (1) defines the one-step-ahead prediction error which is serially uncorrelated by assumption. Therefore, robust methods based on either heteroskedasticity and autocorrelation robust (HAC) standard errors (see Jordà, 2005) or adding additional lags of $w_{t}$ are typically used in practice (see Montiel Olea and Plagborg-Møller, 2021). Another solution, proposed by Lusompa (2023), involves adding estimates of $\{\hat{\varepsilon}_{t+1},\dots,\hat{\varepsilon}_{t+h-1}\}$ (or using a transformation of the left-hand side variable in the projections) to the $h$-step-ahead regressions to obtain valid inference.
All these techniques, however, are based on treating each of the forecasting equations as separate estimation problems. [^8] One exception that takes a Bayesian stance is Tanaka (2020) who relies on a similar representation as we do, and uses it as an estimation device to elicit a joint prior on $\bm{b}$. This has implications for the efficiency of the estimators, since $b_{i}$ and $b_{j}$ are treated independently of each other for each $i\neq j$. Moreover, impulse responses in dynamic models typically feature patterns such as persistence and smoothness, implying that an independent treatment of each horizon complicates incorporating potentially available prior knowledge about the shape of the IRF.
@@
[…]
@@
While the forecast errors of the AR process are uncorrelated, the residuals of the LPs are correlated across horizons because they represent multi-step forecast errors. Our approach takes into account the correlations across horizons — Equation (2) motivates our Bayesian approach to estimate $\bm{b}$, taking into account the correlation structure between the shocks, determined by $\tilde{\bm{Q}}$. We achieve this by introducing a pseudo-likelihood (see also Ferreira et al., 2025, who, however, do not jointly estimate the entire system of equations and instead rely on an ex-post adjustment of the posterior variance of impulse responses), which we build as follows.
@@
Second, at a fundamental level, why is Equation (5) a pseudo-likelihood and why don’t we write down the full likelihood function? The answer is that the set of LPs does not constitute a generative model and as such the prediction error decomposition cannot be applied. To see this, note that a standard prediction error decomposition would force us to use $\mathcal{N}(\bm{y}_{t}|\bm{y}^{t-1},\bm{b},\bm{\Sigma}_{u})$, where the superscript indicates information up to t1t-1. But since $\bm{y}_{t}$ and $\bm{y}_{t-1}$ have elements in common, the prediction error for those elements would be zero, while a standard application of LPs would want to exploit information in the forecast errors for all horizons and time periods. We thus view our pseudo-likelihood as a convenient device that allows us to use the Bayesian toolkit when estimating LPs. As stated in the introduction, relying on a pseudo-likelihood possibly risks model misspecification if the exact (but unknown) likelihood disagrees with the adopted auxiliary likelihood. We return to this issue and discuss how our approach handles specification issues in more detail in Section 3.7.
2.2. Modeling Cross-Horizon Correlations and Dynamic Error Structures
Writing down the likelihood for the entire system takes into account correlations across horizons. To see this, it is useful to decompose $\hat{p}(\bm{y}_{t}|\bm{b},\bm{\Sigma}_{u})$ into: [^10] This decomposition is a prediction-error decomposition, but for one time-period $t$ across all horizons; the problem we just discussed arises when considering a prediction-error decomposition across $t$.
@@ 2.2. Modeling Cross-Horizon Correlations and Dynamic Error Structures
where $\bm{Q}$} is lower triangular with unit diagonal and ${\bm{\Omega}}=\text{diag}({\omega}_{1},\dots,{\omega}_{H})$ are the error variances of $\bm{e}_{t}$, the uncorrelated errors recovered by the Cholesky decomposition. Note that each horizon has a different variance $\omega_{j}$ and hence we have heteroskedasticity across the different forecast horizons. This is in contrast to Equation (3). We view this flexibility as an advantage of our approach, even though it is not strictly needed in this specific application, as a comparison of Equations (2) and the expressions we turn to next show.
@@
[…]
@@
Hence, the $h$-step-ahead LP forecast errors are a function of the independent errors across the preceding h1h-1 horizons as in Equation (1). Our pseudo-likelihood controls for $e_{t}^{(j)}$ for all $j<h$ in the LP for horizon $h$, as these are a function of the parameters of the model and the relevant lags of $w_{t}$. Throughout, we assume that the horizon 0 forecast error $u_{t}^{(0)}$ is serially uncorrelated, a standard assumption in the literature.
To gain a better understanding of the relationship between the different errors, the model can be written in full data notation. When $e_{t+h}^{(i)}=e_{t+h}^{(j)}=e_{t+h}$, we obtain: [^11] A variant capable of strictly enforcing this is to use a state space representation and treat the one-step-ahead prediction errors as unobserved states. We provide a brief discussion in Appendix A.
@@
where the $t$th row of $\bm{U}$ is $\bm{u}_{t}^{\prime}$. The $j$th (for $j=h+1$) column of $\bm{U}$, denoted $\bm{U}_{\bullet j}$, is:
@@
which demonstrates the MA(h)(h) structure of the reduced form errors at horizon $h$ (see, e.g., Chan, 2013, for a discussion in a Bayesian context). It is worth noting that we use this recursive structure merely for expository purposes. In what follows, we will consider full-system estimation based on the unrestricted covariance matrix $\bm{\Sigma}_{u}$, which nests the MA(h)(h) structure but allows for more general correlation structures.
@@ 2.3. General seemingly unrelated linear projections
We start by generalizing the ideas laid out in the previous section. Our goal is to estimate the dynamic response of a target variable $w_{t}$ to a change in a scalar time series $x_{t}$ conditional on a possibly large panel of additional variables stored in $\bm{r}_{t}$ and $\bm{s}_{t}$. [^12] This framework in principle enables us to compute impulse responses for any number of variables, by replacing $w_{t}$ in Equation (7) with a vector of dependent variables $\bm{w}_{t}=(w_{1t},\ldots,w_{Mt})^{\prime}$. These are of dimension $n_{r}$ and $n_{s}$, respectively. The vectors $\bm{r}_{t}$ and $\bm{s}_{t}$ are used to distinguish between predetermined and simultaneously determined covariates. Let $\bm{d}_{t}=(w_{t},x_{t},\bm{r}_{t}^{\prime},\bm{s}_{t}^{\prime})^{\prime}$ and define $\bm{z}_{t}=(\bm{r}_{t}^{\prime},\bm{d}_{t-1}^{\prime},\ldots,\bm{d}_{t-P}^{\prime})^{\prime}$ as a vector of contemporaneous and lagged controls with associated $k\times 1$ parameter vector $\bm{\gamma}_{h}$ where $k=n_{r}+P(n_{r}+n_{s})$.
@@
[…]